OpenStudy (anonymous):

Compute lim x->0 (4x/sin(3x)+x/cos(3x))

myininaya (myininaya):

\[\lim_{x \rightarrow 0}(\frac{4x}{\sin(3x)}+\frac{x}{\cos(3x)}=\frac{4}{3}\lim_{x \rightarrow 0}\frac{3x}{\sin(3x)}+\lim_{x \rightarrow 0}\frac{x}{\cos(3x)}\] Recall \[\lim_{u \rightarrow 0}\frac{u}{\sin(u)}=1\]

myininaya (myininaya):

the other one is just direct sub.

OpenStudy (anonymous):

i didnt quite understand what you just showed =/

myininaya (myininaya):

Nothing?

OpenStudy (anonymous):

okay so what did you do their when you put 4/3

myininaya (myininaya):

4/3 * 3 =4 I just multiplied the first fraction by 3/3

OpenStudy (anonymous):

which became 12x/sin(9x)?

myininaya (myininaya):

oh no

myininaya (myininaya):

3sin(3x) does not equal sin(9x)

OpenStudy (anonymous):

oh ur right lol sorry about that

OpenStudy (anonymous):

whats the purpose of multiplying by 3/3?

myininaya (myininaya):

\[\lim_{u \rightarrow 0}\frac{u}{\sin(u)}=1\] Do you recall this?

OpenStudy (anonymous):

i dont

myininaya (myininaya):

so this means we can say \[\lim_{x \rightarrow 0}\frac{3x}{\sin(3x)}=1\]

OpenStudy (anonymous):

alright

myininaya (myininaya):

so I wanted a 3 up there in our problem

myininaya (myininaya):

but if i put a 3 on top I have to put a 3 on bottom

OpenStudy (anonymous):

yes

myininaya (myininaya):

so that is why I multiplied by 3/3

OpenStudy (anonymous):

i get it so the first part became 1

myininaya (myininaya):

yes 4/3 times 1 is the first part

myininaya (myininaya):

not the second part we just plug in 0

myininaya (myininaya):

now*

myininaya (myininaya):

is everything ok? you got it?

OpenStudy (anonymous):

im writing it down

myininaya (myininaya):

ok

OpenStudy (anonymous):

so then it would be 0/cos(3(0))

myininaya (myininaya):

0/1=0

myininaya (myininaya):

yes

myininaya (myininaya):

so we have 4/3 from the first part+ 0 from the second part =4/3

OpenStudy (anonymous):

oh got it thanks!

myininaya (myininaya):

:)

myininaya (myininaya):

Spikey I appreciate that you want to learn this stuff. Thanks for engaging in the learning process. That is really cool dude.

OpenStudy (anonymous):

lol thank you :) i take AP calc

OpenStudy (anonymous):

in highschool

myininaya (myininaya):

Go you! :)

myininaya (myininaya):

It is really cool to have that attitude. You will go far.

OpenStudy (anonymous):

theres just some stuff that i dont really comprehend my teacher doesnt teach some stuff clearly

OpenStudy (anonymous):

lol yeah i never give up. :)

OpenStudy (anonymous):

always trying to learn new things.

myininaya (myininaya):

Good luck to you and peace out!

OpenStudy (anonymous):

lol bye